Subject: 9/10/01 APO 3.5m User's Committee Minutes

From: strauss@astro.Princeton.EDU

Submitted: Mon, 17 Sep 2001 15:39:05 -0400 (EDT)

Message number: 525 (previous: 524, next: 526 up: Index)

  Apache Point Observatory 3.5m User's Committee Meeting
		September 10, 2001

Attending: Jon Holtzman, Bruce Balick, Rene Walterbos, Mike Shull,
Bruce Gillespie, Ed Turner, Michael Strauss, Lew Hobbs, Alan Uomoto

Not attending: Chris Stubbs

************Status of Baffling Project********************************
 A detailed report from the consulting company has now arrived, with a
detailed model of scattering in the telescope.   This will become a
front-burner project as soon as the DIS upgrade is done (see below).  
In particular, John Davis is putting work on the baffling at high priority. 
  People are fairly confident that the model from the consulting
company is quite good; it qualitatively describes the pinhole models
quite well.  
  We've already shrouded the tertiary column as recommended by the
consulting company; tests to see if this makes a difference (a series
of pinhole images, as well as other tests suggested by Jon Holtzman)
will happen soon. 
  The consulting company made a series of recommendations, from easy
to hard to implement, with an estimate of the effect of each.  To
understand how to proceed, we need some quantitative science goals
here.  The ultimate desiratum is how well we can flat-field; right
now, we can do so only to roughly 10% over the field of SPICAM.  Holtzman
suggests a goal of 1% flatfielding in dark time (i.e., the ability to
do reproducible photometry of a given object over the field to 1%),
although it is not clear over what field that goal should be set: the
5 arcmin of SPICAM, or the full 30 arcmin that the telescope delivers? 
  There is some confusion about exactly what the effects of the
scattered light actually are on flat fielding and photometry.  Turner
points out that in his quasar monitoring program, he has been
successful in doing 1% photometry using a SPICAM superflat derived
from many nights of data.  It would be interesting to hear other
peoples' experience in trying to do accurate photometry with the
telescope. 

**********************The DIS Upgrade*********************************
  The DIS upgrade remains on track.  The Cryotiger is in hand; on
engineering chip is in one of the dewars already, being used to check
out the software and electronic noise of the system.  Then the
observing grade chip will be installed and characterized.  So we hope
that this will be accomplished in early October. 

  There is a possibility that the upgrade will not be accomplished by
the end of the October bright run; if the preparation is not complete
by the beginning of the dark run, the whole upgrade could be put off
for the November dark run.  Alternatively, it might happen that the
work on the instrument will be started, but not be complete by the end
of the October bright run, which of course would not allow observation
with DIS during the October dark run. In this case, DIS observers
should be ready with possible backup projects using other instruments.
We will also consider switching observing time to other highly ranked
proposals from the same institution.

**********************Colorado near-IR instrument (NICFPS)*********** 
Ball Aerospace will contribute 200 hours of in-kind design work,
reviewing the plans and designs of the CU group.  They are about to
put in the order for a Hawaii 1KX1K chip.  They have a grad student
full-time, and a half-time postdoc working on this.

***********************The New Top End********************************
It is becoming increasingly clear that our original approach of doing
the design in two stages (frame first, mounting of secondary and
controlling its motion) is not going to work because there's a lot of
engineering "coupling" in the designs of the two stages.  So we're going to have
to do both aspects of the project together, on the timescale of next year.

**********************Status of DIMM (seeing monitor)****************
  DIMM hasn't been in regular service since March.  There is an SDSS
plan to move it to a permanent mounting at the end of the Sloan pier.
Its old location near the 1-meter telescope suffered from bad local 
air, and the new location has been shown to be in "free" air and 
correlates well with what the 3.5-m sees.

**************The Just-Completed Summer Shutdown*********************
The shutdown, which was completed a week or so ago, was quite
successful, perhaps the most successful we've had yet.  The appendix
to these minutes lists the major accomplishments.  Essentially
everything planned to be done was done (and was finished a few days
early).  Moreover, the monsoon coincided almost perfectly with the
shutdown.  Shack-Hartmann tests show that we have better collimation
than ever before, and we've seen superb seeing since then ( 0.5").
The intrinsic delivered image quality from the optics is 0.2" in the
visible.  Finally, the small cracks in the primary from a few years
ago have *not* grown in size; this is reassuring news, as this means
they are not likely to grow further in the future. 

*********************New Observing Specialist***********************
  We have a new 1/2 time 3.5m observing specialist, John Barentine
shared with the APO 1m; he has started training, and is quickly
coming up to speed.  Welcome!

  We have considering having the observing specialist contact the
observer of each night directly, to make sure they are all ready to go
(e.g., if they need to do calibrations in the afternoon).

Last month's minutes are approved. 

Next meeting 11:30 AM East Coast Time, Monday, October 8. 

**********Appendix: Accomplishments of the shutdown. *****************

1. Both enclosure rotation motors were repaired which required
replacing bearings and repairing one endplate that was damaged by
the brake assembly over the years.

2. The primary mirror was washed

3. The tertiary mirror was re-aluminized

4. UBVR and I filters were added to the guider

5. The rotator wrap was changed to allow +/- 360 deg rotation and
make instrument changes easier

6. The guider heat exchanger was serviced and changed to circulate
outside air versus air trapped in the guider box

7. The mirror support system was serviced  fixing air leaks,
improved switch over functions to the backup system, checked random
bellofram pistons for damage, hardening and corrosion, calibrated
the load cells and replaced  transverse amplifier/load cell
combination, replaced 1 servo board and added resistors to improve
tuneability

8. Inspected primary mirror

9. Installed Printed circuit board for PCB3 on tertiary

10. Repaired tertiary rotation "at slot" sensor

11. Adjusted Echelle focus

12. Adjusted primary mirror vent tubes to provide improved air circulation

13. Replaced casters on GFP instrument cart

14. Serviced telescope azimuth bearing

15. Serviced telescope azimuth moat seal

16. Installed new design for tertiary wiffle which eliminates 1
known failure mode, improves maintainability, and installed larger
flex pivots providing a slightly stiffer design

17. Evacuated echelle upper tank, Grim2, Dis  and Spicam

18. Rearranged Instrument umbilicals and cleaned up attach points

19. Installed telemetry on elevation and rotator drives and azimuth LVDT

20. Replaced Tycho

21. Installed new TCC version and made some MC software changes

22. Relocated Shack-Hartmann instrument to port MC4


APO APO APO APO APO  Apache Point Observatory 3.5m  APO APO APO
APO
APO  This is message 525 in the apo35-general archive. You can find
APO  the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO  To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO  To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO